# HyConEx: Hypernetwork classifier with counterfactual explanations

### v1

@@ Abstract
In recent years, there has been a growing interest in explainable AI methods. We want not only to make accurate predictions using sophisticated neural networks but also to understand what the model’s decision is based on. One of the fundamental levels of interpretability is to provide counterfactual examples explaining the rationale behind the decision and identifying which features, and to what extent, must be modified to alter the model’s outcome. To address these requirements, we introduce HyConEx, a classification model based on deep hypernetworks specifically designed for tabular data. Owing to its unique architecture, HyConEx not only provides class predictions but also delivers local interpretations for individual data samples in the form of counterfactual examples that steer a given sample toward an alternative class. While many explainable methods generated counterfactuals for external models, there have been no interpretable classifiers simultaneously producing counterfactual samples so far. HyConEx achieves competitive performance on several metrics assessing classification accuracy and fulfilling the criteria of a proper counterfactual attack. This makes HyConEx a distinctive deep learning model, which combines predictions and explainers as an all-in-one neural network. The code is available at https://github.com/gmum/HyConEx.
Keywords:
Explainable AI Counterfactual examples Hypernetworks Invertible Normalizing Flows Tabular Data.
Figure 1: Comparison of HyConEx with related explanability methods. Feature attribution methods, like LIME or SHAP, deliver static explanations supporting only current decision. External methods for generating counterfactual examples, eg. PPCEF or CEGP, typically perform time-consuming optimization for every input data point. HyConEx enclose predictions and counterfactuals in a single method, which makes it extremely efficient method that delivers dynamic explanation.
@@ 1 Introduction
Explanations can be delivered at several levels. For instance, banks due to legal reasons need to explain refusing or granting a loan. These explanations define the importance of input features, which are crucial for making a given decision. On the other hand, a bank customer would like to know what she/he should do to change the bank’s decision [17]. This topic is related to counterfactual explanations informing which features, and to what extent, have to be modified to obtain a desired model output, e.g. getting a loan. Feature importance methods deliver static information about classification (factual decision). In contrast, counterfactual examples are a type of causal reasoning and provide better insight into the dynamic of the classification rule, see Figure 1.
While there exist highly accurate machine learning classifiers for tabular data [15, 16, 19, 21], as well as external methods for generating counterfactual explanations  [26, 33, 38], no existing model combines these two objectives. In this paper, we make a step towards interpretable deep learning and combine the classification power of neural networks with the dynamic interpretability of counterfactual examples. We introduce HyConEx, a type of hypernetwork, which for every input returns a local classifier equipped with counterfactual vectors targeting all other classes. In this way, HyConEx not only returns a highly accurate decision but also suggests possible options of how to change that decision. Enclosing these two features in one model allows us to speed-up the inference and avoid time-consuming optimization of counterfactuals generation, which is common for external counterfactual explainers [38], see Figure 1. To the best of our knowledge, HyConEx is the first deep learning approach having these two features enclosed in one model.
Our experiments demonstrate that HyConEx achieves accuracy competitive to the state-of-the-art methods (Table 1). Moreover, it produces counterfactual explanations for all classes of comparable quality to external counterfactual explainers (Tables 2, 3 and 4). Since counterfactuals are generated jointly with classification scores in a single forward pass, it does not introduce any additional computational overhead, which makes it a preferred choice over external methods that require time-consuming optimization.
@@ 2.1 Classical Interpretable Models
Tabular data are ubiquitous in domains such as finance, medicine, and fraud detection. Traditional models such as logistic regression and decision trees have long been favored for their inherent transparency. Logistic regression [20], for instance, has been extensively studied due to its linear formulation that directly relates feature coefficients to prediction outcomes. Similarly, linear Support Vector Machines (SVMs) [11] offer interpretability through the explicit characterization of support vectors and decision boundaries. Decision trees provide a natural form of explanation by decomposing decisions into a sequence of simple, human-readable rules. However, while tree-based ensemble methods (for example, boosted trees such as XGBoost [10], LightGBM [22], and CatBoost [15]) achieve high predictive performance in tabular data, the aggregated nature of these models often obscures the interpretability of individual decisions, necessitating the use of additional explanation techniques.
@@ 2.2 Post-Hoc and Counterfactual Explanations
To mitigate the opacity of high-performing black-box models, several post-hoc explanation techniques have been proposed. Methods such as LIME (Local Interpretable Model-agnostic Explanations) [32] and SHAP (SHapley Additive exPlanations) [25] approximate complex models locally with simpler, interpretable surrogates that assign importance scores to individual features. These approaches have proven effective in providing insights into the local behavior of classifiers, albeit sometimes at the cost of stability and global consistency.
Counterfactual explanation methods have emerged as a complementary paradigm by identifying minimal changes in input features that would alter a model prediction, thus offering actionable insights into how a different outcome might be achieved. These methods can be broadly categorized into endogenous explainers, which generate counterfactuals by selecting or recombining feature values from the existing dataset, and exogenous explainers, which produce counterfactual examples through interpolation or random data generation without the strict constraint of naturally occurring feature values [17]. Early work by Wachter et al. [37] laid the foundation for exogenous methods, and subsequent approaches have enhanced plausibility by imposing constraints such as convex density restrictions proposed by Artelt and Hammer [5], integrating density estimation via normalizing flows as in PPCEF [38], or guiding the search for interpretable counterfactuals with class prototypes as in CEGP [35]. On the endogenous side, the Contrastive Explanations Method (CEM) [12] emphasizes the importance of both pertinent positives and negatives to justify a prediction, an approach further refined by FACE [29], which identifies feasible and actionable modification paths that respect inherent data and decision constraints, and by the Case-Based Counterfactual Explainer (CBCE) [23] that forms explanation cases by pairing similar instances with contrasting outcomes. Although these techniques vary in their mechanisms, they commonly require additional post hoc optimization separate from the classifier.
@@ 2.3 Deep Interpretable Models
Deep interpretable models have emerged as a promising avenue to reconcile high predictive performance with intrinsic explainability in tabular data. The early work of Alvarez-Melis and Jaakkola [3] pioneered the idea of self-explaining neural networks that enforce local linearity and stability, laying the groundwork for models that can be interpretable by design. Based on these concepts, attention-based architectures such as TabNet [4] have demonstrated that sequential attention mechanisms can dynamically select the most relevant features during decision-making, offering built-in interpretability alongside competitive accuracy. Later, architectures like the Interpretable Mesomorphic Neural Networks (IMN) [21] leverage deep hypernetworks to generate instance-specific linear models, providing direct feature attributions without compromising the expressive power of deep networks. Complementary approaches include Neural Additive Models (NAMs) [2], which combine the flexibility of deep learning with the transparency of generalized additive models through per-feature shape functions, and Deep Abstract Networks (DANet) [9], which utilize specialized abstract layers to hierarchically group and transform features for clearer semantics.
@@
3 HyConEx model
This section describes HyConEx, which is a deep learning classifier equipped with counterfactual explanations. Since HyConEx is an all-in-one neural network, predictions and counterfactuals are returned in a single forward pass. First, we introduce the basic concept of counterfactual explanations. Next, we present two main components of HyConEx: hypernetwork [1] and invertible normalizing flows [31]. Finally, we describe HyConEx model in detail.
@@ 3.1 Counterfactual explanations
To describe a basic approach to generating counterfactual examples, let ${f:\mathbb{R}^{D}\to\mathbb{R}^{K}}$ be a multi-class classifier trained on a given dataset $\{(x_{i},y_{i})\}_{i=1}^{N}$, where $D$ denotes the number of input features, $K$ is the number of classes and $N$ is the number of training data. We assume that the classifier $f$ returns a class label $y$ for a data point $x$. The question is: how to modify $x$ to obtain an alternative class $y^{\prime}$? Since we are interested in local decisions, we want to make as small changes to $x$ as possible. Formally, counterfactual example $x^{\prime}$ can be defined as a solution to the optimization problem [37]:
$\mathrm{argmin}_{x^{\prime}\in\mathbb{R}^{D}}\mathcal{L}(f(x^{\prime}),y^{\prime})+C\cdot d(x,x^{\prime}),$
@@
Contemporary methods satisfy many additional properties, such as sparsity, actionability, or diversity, which increase their interpretation capabilities [36]. In this paper, we are especially interested in constructing plausible counterfactuals [17]. A counterfactual is plausible if it is located within the high-density region of data. This feature distinguishes counterfactuals from adversarial examples. Plausibility is typically assessed via a straightforward k-neighborhood analysis of the counterfactual relative to the original dataset [6]. Recently, PPCEF [38] has employed invertible normalizing flows to evaluate plausibility in tabular data contexts. Our study adopts a similar approach by integrating invertible normalizing flows directly into the classification model.
@@ 3.2 Hypernetwork classifiers
Hypernetworks, introduced in [1], are defined as neural models that generate weights for a separate target network that solves a specific task. In this seminal work, the authors aim to reduce the number of trainable parameters by designing a hypernetwork with fewer parameters than the target network. The hypernetworks were developed and applied in several areas including generative modeling, continual learning [27], image representations [24], and many others[34].
The recently introduced Interpretable Mesomorphic Networks (IMN) [21] uses a hypernetwork to construct an individual local linear classifier for every tabular data point. Due to the linearity of the classifier, IMN delivers the local importance of the input features (determined by the linear model). Although IMN uses local linear models, it produces global non-linear decision boundary in the whole data space.
In this approach, the hypernetwork $H(\cdot;\theta):\mathbb{R}^{D}\to\mathbb{R}^{K\times(D+1)}$ takes a data point $x\in\mathbb{R}^{D}$ and outputs the local linear classifier for $K$-class problem given by the weight matrix $W=H(x;\theta)$. The $k$-th row of $W$ determines the logit of the $k$-th class:
$z(x;W)_{k}=\sum_{d=1}^{D}W_{k,d}x_{d}+W_{k,0}.$
In consequence, the probability that $x$ belongs to $k$-th class equals:
$f(x;W)_{k}=\frac{\exp({z(x;W)_{k}})}{\sum_{j=1}^{K}\exp({z(x;W)_{j}})}.$
The hypernetwork is instantiated by Tabular ResNet [21].
Figure 2: Illustration of all elements of our cost function: (a) Cross-entropy is applied only to the input example similar to IMN [21]. As we can see, the normal vector $w$ is turned into the target class, but the translation of $x^{\prime}=x-w$ does not lead to the change of class label; (b) We additionally apply the cross-entropy to the counterfactual example. In this case, we produce strong modification, which does not lead to in-distribution samples; (c) We supply the previous loss with the proximity between input and counterfactual. We observe that the model gives the smallest modification, which changes the class label. (d) Complete loss with plausibility term. The final HyConEx enforces $x^{\prime}$ to lie in the manifold of the target class.
@@ 3.3 Invertible Normalizing Flows
To generate plausible counterfactuals, we use invertible normalizing flows (INFs) to model the density of each class [31].
@@
INFs transform a latent variable $z$ with a known prior distribution $p(z)$ into an observed space variable $x$ with an unknown distribution. This transformation is modeled by a series of invertible (parametric) functions:
$x=F(x)=f_{K}\circ\ldots\circ f_{1}(z).$
@@
$\log p_{F}(x|y)=\log p(z)-\sum^{K}_{k=1}\log\bigg{|}det\frac{\partial f_{k}}{\partial z_{k-1}}\bigg{|},$
where $z=f^{-1}_{1}\circ\ldots\circ f^{-1}_{K}(x,y)$ is a result of the invertible mapping. A significant challenge in INFs is the selection of appropriate invertible functions, denoted as $f_{K},\dots,f_{1}$. The existing literature presents a variety of solutions to address this issue, with notable methodologies including NICE [13], RealNVP [14], and MAF [28].
3.4 HyConEx architecture
Our architecture comprises two key elements: hypernetwork classifier and invertible normalizing flows, see Figure 3. By defining a proper loss function and designing dedicated training procedure, the generated classifier delivers predictions and the associated counterfactual explanations.
Figure 3: Illustration of HyConEx architecture. The hypernetwork takes every input together with the estimated data density and returns: (1) local decision boundary, and (2) counterfactual examples for all alternative classes. In consequence, we obtain predictions and counterfactuals in a single forward pass. We emphasize that HyConEx uses a single hypernetwork, which was duplicated here into $n$ models for illustration only.

### v2

@@ Abstract
In recent years, there has been a growing interest in explainable AI methods. In addition to making accurate predictions, we also want to understand what the model’s decision is based on. One of the fundamental levels of interpretability is to provide counterfactual examples explaining the rationale behind the decision and identifying which features, and to what extent, must be modified to alter the model’s outcome. To address these requirements, we introduce HyConEx, a classification model based on deep hypernetworks specifically designed for tabular data. Owing to its unique architecture, HyConEx not only provides class predictions but also delivers local interpretations for individual data samples in the form of counterfactual examples that steer a given sample toward an alternative class. While many explainable methods generate counterfactuals for external models, there have been no interpretable classifiers simultaneously producing counterfactual samples so far. HyConEx achieves competitive performance on several metrics assessing classification accuracy and fulfilling the criteria of a proper counterfactual attack. This makes HyConEx a distinctive deep learning model, which combines predictions and explainers as an all-in-one neural network. The code is available at https://github.com/gmum/HyConEx. This work has been published in Neurocomputing and is available at: https://doi.org/10.1016/j.neucom.2026.132748.
Keywords Explainable AI $\cdot$ Counterfactual examples $\cdot$ Hypernetworks $\cdot$ Invertible Normalizing Flows $\cdot$ Tabular Data.
@@ 1 Introduction
Explanations can be delivered at several levels. For instance, banks, due to legal reasons, need to explain why they refuse or grant a loan. These explanations define the importance of input features, which are crucial for making a given decision. On the other hand, a bank customer would like to know what she/he should do to change the bank’s decision [5]. This topic is related to counterfactual explanations informing which features, and to what extent, have to be modified to obtain a desired model output, e.g. getting a loan. Feature importance methods deliver static information about classification (factual decision). In contrast, counterfactual examples are a type of causal reasoning and provide better insight into the dynamic of the classification rule.
While there exist highly accurate machine learning classifiers for tabular data [3, 6, 1, 7], as well as external methods for generating counterfactual explanations  [8, 9, 10], no existing model combines these two objectives. In this paper, we make a step towards interpretable deep learning and combine the classification power of neural networks with the dynamic interpretability of counterfactual examples. We introduce HyConEx, a type of hypernetwork, which for every input returns a local classifier equipped with counterfactual vectors targeting all other classes. In this way, HyConEx not only returns a highly accurate decision but also suggests possible options of how to change that decision. Unlike existing interpretable classifiers that provide feature importance scores explaining why a decision was made, HyConEx delivers counterfactual examples that answer how to change that decision through actionable guidance generated in a single forward pass. Enclosing classification and counterfactual generation in one model allows us to speed-up the inference and avoid time-consuming optimization of counterfactuals generation, which is common for external counterfactual explainers [10]. To the best of our knowledge, HyConEx is the first deep learning approach having these two features enclosed in one model.
Our experiments demonstrate that HyConEx achieves accuracy competitive to the state-of-the-art methods (Table 3). Moreover, it produces counterfactual explanations for all classes of comparable quality to external counterfactual explainers (Tables 5, 6 and 7). Since counterfactuals are generated jointly with classification scores in a single forward pass, it does not introduce any additional computational overhead, which makes it a preferred choice over external methods that require time-consuming optimization.
@@ 2.1 Classical Interpretable Models
Tabular data are ubiquitous in domains such as finance, medicine, and fraud detection. Traditional models such as logistic regression and decision trees have long been favored for their inherent transparency. Logistic regression [11], for instance, has been extensively studied due to its linear formulation that directly relates feature coefficients to prediction outcomes. Similarly, linear Support Vector Machines (SVMs) [12] offer interpretability through the explicit characterization of support vectors and decision boundaries. Decision trees provide a natural form of explanation by decomposing decisions into a sequence of simple, human-readable rules. However, while tree-based ensemble methods (for example, boosted trees such as XGBoost [4] and CatBoost [3]) achieve high predictive performance in tabular data, the aggregated nature of these models often obscures the interpretability of individual decisions, necessitating the use of additional explanation techniques.
@@ 2.2 Post-Hoc and Counterfactual Explanations
To mitigate the opacity of high-performing black-box models, several post-hoc explanation techniques have been proposed [13, 14, 15]. Methods such as Local Interpretable Model-agnostic Explanations (LIME) [16] and SHapley Additive exPlanations (SHAP) [17] approximate complex models locally with simpler, interpretable surrogates that assign importance scores to individual features. These approaches have proven effective in providing insights into the local behavior of classifiers, albeit sometimes at the cost of stability and global consistency.
Counterfactual explanation methods have emerged as a complementary paradigm by identifying minimal changes in input features that would alter a model prediction, thus offering actionable insights into how a different outcome might be achieved. These methods can be broadly categorized into endogenous explainers, which generate counterfactuals by selecting or recombining feature values from the existing dataset, and exogenous explainers, which produce counterfactual examples through interpolation or random data generation without the strict constraint of naturally occurring feature values [5]. Early work by Wachter et al. [18] laid the foundation for exogenous methods, and subsequent approaches have enhanced plausibility by imposing constraints such as convex density restrictions proposed by Artelt and Hammer [19], integrating density estimation via normalizing flows as in Probabilistically Plausibile Counterfactual Explanations using Normalizing Flows (PPCEF) [10], or guiding the search for interpretable counterfactuals with class prototypes as in Counterfactual Explanations Guided by Prototypes (CEGP) [20]. On the endogenous side, the Contrastive Explanations Method (CEM) [21] emphasizes the importance of both pertinent positives and negatives to justify a prediction, an approach further refined by Feasible and Actionable Counterfactual Explanations (FACE) [22], which identifies feasible and actionable modification paths that respect inherent data and decision constraints, and by the Case-Based Counterfactual Explainer (CBCE) [23] that forms explanation cases by pairing similar instances with contrasting outcomes. Although these techniques vary in their mechanisms, they commonly require additional post hoc optimization separate from the classifier.
@@ 2.3 Deep Interpretable Models
Deep interpretable models have emerged as a promising avenue to reconcile high predictive performance with intrinsic explainability in tabular data. The early work of Alvarez-Melis and Jaakkola [24] pioneered the idea of self‐explaining neural networks that enforce local linearity and stability, laying the groundwork for models that can be interpretable by design. Based on these concepts, attention‐based architectures such as TabNet [25] have demonstrated that sequential attention mechanisms can dynamically select the most relevant features during decision-making, offering built-in interpretability alongside competitive accuracy. Later, architectures like the Interpretable Mesomorphic Neural Networks (IMN) [7] leverage deep hypernetworks to generate instance‐specific linear models, providing direct feature attributions without compromising the expressive power of deep networks. Complementary approaches include Neural Additive Models (NAMs) [26], which combine the flexibility of deep learning with the transparency of generalized additive models through per-feature shape functions, and Deep Abstract Networks (DANet) [27], which utilize specialized abstract layers to hierarchically group and transform features for clearer semantics.
@@
3 Preliminaries
Table 1: Main notation used throughout the paper.
Symbol Description
Dims $D$ Number of input features
$K$ Number of classes
NN Number of training samples
Variables $x\in\mathbb{R}^{D}$ Input data point
$x^{\prime}\in\mathbb{R}^{D}$ Counterfactual example
$y\in\{1,\ldots,K\}$ True class label
$k\in\{1,\ldots,K\}$ Predicted class label
$m\in\{1,\ldots,K\}$ Target class for counterfactual
$W\in\mathbb{R}^{K\times(D+1)}$ Weight matrix output by hypernetwork
$W_{m}$ $m$-th row of weight matrix $W$
$z(x;W)\in\mathbb{R}^{K}$ Logits vector for $x$ determined by $W$
$z(x;W)_{k}$ Logit of the $k$-th class
$\theta$ Parameters of hypernetwork $H$
$\phi$ Parameters of normalizing flow $F$
Functions & Notations $f:\mathbb{R}^{D}\to\mathbb{R}^{K}$ Classifier function
$H(\cdot;\theta):\mathbb{R}^{D}\to\mathbb{R}^{K\times(D+1)}$ Hypernetwork with parameters $\theta$
$F(\cdot|\cdot;\phi)$ Conditional normalizing flow with parameters $\phi$
$\mathcal{L}_{CE}$ Cross-entropy loss
$\mathcal{L}_{conEx}$ Counterfactual loss
$\mathcal{L}_{F}$ Flow plausibility loss
$d(\cdot,\cdot)$ Distance measure
$p_{F}(x|y)$ Conditional probability density
$\alpha_{1},\alpha_{2},\alpha_{3}$ Loss regularization hyperparameters
$\delta$ Density threshold for plausibility
To describe HyConEx, we first introduce the basic concept of counterfactual explanations. Next, we present two main components of HyConEx: hypernetwork [28] and invertible normalizing flows [29]. The key symbols used throughout the paper are summarized in Table 1.
Figure 1: Comparison of HyConEx with related explanability methods. Feature attribution methods, like LIME or SHAP, deliver static explanations supporting only the current decision. External methods for generating counterfactual examples, e.g. PPCEF or CEGP, typically perform time-consuming optimization for every input data point. HyConEx encloses predictions and counterfactuals in a single forward pass, which makes it an extremely efficient method that delivers a dynamic explanation.
@@ 3.1 Counterfactual explanations
To describe a basic approach to generating counterfactual examples, let $f:\mathbb{R}^{D}\to\mathbb{R}^{K}$} be a multi-class classifier trained on a given dataset $\{(x_{i},y_{i})\}_{i=1}^{N}$, where $D$ denotes the number of input features, $K$ is the number of classes and $N$ is the number of training data. We assume that the classifier $f$ returns a class label $k$ for a data point $x$. The question is: how to modify $x$ to obtain an alternative class $m$? Since we are interested in local decisions, we want to make as small changes to $x$ as possible. Formally, counterfactual example $x^{\prime}$ can be defined as a solution to the optimization problem [18]:
$\mathrm{argmin}_{x^{\prime}\in\mathbb{R}^{D}}\mathcal{L}(f(x^{\prime}),m)+C\cdot d(x,x^{\prime}),$ (1)
@@
Contemporary methods satisfy many additional properties, such as sparsity, actionability, or diversity, which increase their interpretation capabilities [30]. In this paper, we are especially interested in constructing plausible counterfactuals [5]. A counterfactual is plausible if it is located within the high-density region of data. This feature distinguishes counterfactuals from adversarial examples. Plausibility is typically assessed via a straightforward k-neighborhood analysis of the counterfactual relative to the original dataset [31]. Recently, PPCEF [10] has employed invertible normalizing flows to evaluate plausibility in tabular data contexts. Our study adopts a similar approach by integrating invertible normalizing flows directly into the classification model. A brief comparison of HyConEx with other methods delivering counterfactuals and decision explanations is presented in Figure 1.
@@ 3.2 Hypernetwork classifiers
Hypernetworks, introduced in [28], are defined as neural models that generate weights for a separate target network that solves a specific task. The recently introduced IMN [7] uses a hypernetwork to construct an individual local linear classifier for every tabular data point. Although IMN uses local linear models, it produces a global non-linear decision boundary in the whole data space.
In this approach, the hypernetwork $H(\cdot;\theta):\mathbb{R}^{D}\to\mathbb{R}^{K\times(D+1)}$}, parameterized by weights $\theta$, takes a data point $x\in\mathbb{R}^{D}$ and outputs the local linear classifier for $K$-class problem given by the weight matrix $W=H(x;\theta)$. The $k$-th row of $W$ (denoted by $W_{k}$) determines the logit of the $k$-th class as follows:
$z(x;W)_{k}=\sum_{d=1}^{D}W_{k,d}x_{d}+W_{k,0},$ (2)
where $z(x;W)_{k}$ is the $k$-th coordinate of the vector $z(x;W)$. In consequence, the probability that $x$ belongs to $k$-th class is parametrized by applying a softmax function to the logits $z(x,W)$:
$f(x;W)_{k}=\frac{\exp({z(x;W)_{k}})}{\sum_{j=1}^{K}\exp({z(x;W)_{j}})}.$ (3)
Since prediction for every data point is determined by possibly different linear model, the overall decision boundary on the data space can be highly non-linear. The hypernetwork is instantiated by Tabular ResNet [7].
(a) Pure cross-entropy for the input sample: no class transition.
(b) Addition of the cross-entropy loss for the counterfact: class transition, out-of-distribution sample, long distance between the input and the counterfact.
(c) Addition of the proximity loss: minimization of the distance between the input and the counterfact, class transition, out-of-distribution counterfact.
(d) Addition of the plausibility loss: ensuring the in-distributed counterfactual sample.
Figure 2: Illustration of the impact of all components of our cost function: (a) Cross-entropy is applied only to the input example similar to IMN [7]. As we can see, the normal vector $w$ is turned into the target class, but the translation of $x^{\prime}=x-w$ does not lead to the change of class label; (b) We additionally apply the cross-entropy to the counterfactual example. In this case, we produce a strong modification, which does not lead to in-distribution samples; (c) We supply the previous loss with the proximity between input and counterfactual. We observe that the model gives the smallest modification, which changes the class label. (d) Complete loss with plausibility term. The final HyConEx enforces $x^{\prime}$ to lie in the manifold of the target class.
@@ 3.3 Invertible Normalizing Flows
To generate plausible counterfactuals, we need to model the density of data [29]. This will be estimated using invertible normalizing flows (INFs).
@@
INFs transform a latent variable $z$ with a known prior distribution $p(z)$ into an observed space variable $x$ with an unknown distribution. This mapping is achieved through a sequence of invertible parametric transformations:
$x=F(z)=g_{L}\circ\ldots\circ g_{1}(z),$ (4)
where each function $g_{l}$, for $l=1,\ldots,L$, progressively transforms the latent representation toward the observed data space.
@@
$\log p_{F}(x|y)=\log p(z)-\sum^{L}_{l=1}\log\bigg|det\frac{\partial g_{l}}{\partial z_{l-1}}\bigg|,$ (5)
where $z=z_{0}=g^{-1}_{1}\circ\ldots\circ g^{-1}_{L}(x,y)$ is the latent obtained by inverting the flow, and $z_{l}=g_{l}(z_{l-1})$ denotes the intermediate flow variable at layer $l$. A significant challenge in INFs is the selection of appropriate invertible functions, The existing literature presents a variety of solutions to address this issue, with notable methodologies including Non-linear Independent Components Estimation (NICE) [32], Real-Valued Non-Volume Preserving (RealNVP) [33], and Masked Autoregressive Flow (MAF) [34].
4 HyConEx model
This section describes HyConEx, which is a deep learning classifier equipped with counterfactual explanations. Since HyConEx is an all-in-one neural network, predictions and counterfactuals are returned in a single forward pass. Table 1 summarizes the key notation used in our model description.
4.1 Architecture
Our architecture comprises two key elements: hypernetwork classifier and invertible normalizing flows, see Figure 3. By defining a proper loss function and designing a dedicated training procedure, the generated classifier delivers predictions and the associated counterfactual explanations.
Figure 3: Illustration of HyConEx architecture. The hypernetwork takes every input together with the estimated data density and returns: (1) a local decision boundary, and (2) counterfactual examples for all alternative classes. In consequence, we obtain predictions and counterfactuals in a single forward pass.